# Acceleration without photon pair creation

### v1

@@ Abstract
Since quantum physics is more complex, it cannot be derived from classical theories and additional postulates are needed. However, whenever an experiment can be described classically, quantum physics must predict the same outcome. Otherwise, it would be possible to contradict quantum theories with classical experiments. Intuitively, an accelerating observer travelling through a vacuum seems to be a situation that can be analysed classically. It is, therefore, not surprising that many people are puzzled by the Unruh effect, which predicts that the observer encounters photons in a thermal state. This paper shows that the quantised electromagnetic field in a non-inertial frame can be modelled consistently without transforming the vacuum into a thermal state.
@@
Introduction. Despite its doubtless success, quantum physics still makes some unusual predictions that continue to puzzle scientists and raise questions about the validity of some of its aspects. One example is the Unruh effect [1, 2, 3], which predicts that a uniformly accelerating observer would detect a thermal bath of photons where a resting observer would not see any. The creation of photon pairs in the frame of an accelerating observer seems necessary in order to reconcile relativity with quantum electrodynamics [7, 9, 4, 8, 6, 5]; however, the standard explanation of the Unruh effect does not provide much physical intuition into this process, and relies instead on mostly mathematical arguments. Since being proposed in 1976, the Unruh effect has been confirmed by analogue gravity experiments [10, 12, 11, 13, 14, 15]. These, however, are not independent confirmations of photon pair creation in the reference frame of an accelerating observer since they are designed to exhibit a certain expected behavior. In addition, a few non-analogue experiments studied the interaction between accelerating charged particles with the surrounding free radiation field, but they remained inconclusive regarding the question of photon pair creation from the vacuum state [16, 17, 18, 19].
@@
Figure 1: (a) A spacetime diagram illustrating the Doppler effect. Here, a moving observer (Bob) travels at a constant velocity $v_{\rm B}$ (green line) away from a resting observer (Alice). In addition, we assume that Alice sends light signals to Bob at intervals of $\Delta t_{\rm A}$, which he receives at constant time intervals $\Delta t_{\rm B}$. (b) A spacetime diagram illustrating the Unruh effect. Now Bob moves with a time-dependent velocity $v_{\rm B}(t_{\rm A})<c$ (green line) away from Alice. Notice that Bob’s dynamics looks the same in both figures during a short time interval $\Delta t_{\rm B}$ as his trajectory can be approximated by a collection of short straight lines at velocity $v_{\rm B}^{(m)}$ with $m=1,2,3,...$
@@
An effect that is closely related to the Unruh effect, but much better understood, is the relativistic Doppler effect [20, 21, 22]. Suppose a stationary observer (let us call her Alice) sends short light pulses times $\Delta t_{\rm A}$ apart to another observer (let us call him Bob) who is moving at a constant velocity $v_{\rm B}>0$ away from Alice (cf. Fig. 1(a)). From special relativity, we know that the speed of light $c$ is the same in the reference frames of both observers, although they experience space and time differently. Since the time $\Delta t_{\rm B}$ between two subsequent light pulses arriving at Bob’s position is larger than $\Delta t_{\rm A}$, time passes slower in the moving frame. Another way of describing the Doppler effect is to notice that the frequency of monochromatic waves of light is higher in Bob’s reference frame than in Alice’s. Most importantly, if the electromagnetic (EM) field is in its vacuum state in one frame, the same is true in the other. Taking this into account, the particle generation predicted by the Unruh effect seems even more mysterious since the trajectory of an accelerating observer can be approximated by a collection of short time intervals during which the observer moves at a constant speed $v_{\rm B}^{(m)}$ with $m=1,2,3,...$ (cf. Fig. 1(b)).
@@
In this paper, we therefore revisit the validity of the standard conclusions of light propagation in non-inertial reference frames by making use of a recent physically motivated quantisation scheme of the EM field in position space [23, 24]. The main advantage of considering local instead of monochromatic photons as the basic building blocks of the quantised EM field is that this allows us to make use of the principle of relativity and of spatial and time translational symmetries in a more direct way. Moreover, we can avoid the approximation of standing modes in a flat spacetime, which usually accompanies the derivation of the Unruh effect [25]. As we shall see below, our alternative description of the EM field suggests that the vacuum state is the same in all non-inertial reference frames —only zero point energies change. Consequently, the situation in Fig. 1(b) can be modelled consistently without the need for thermal photons.
@@
For simplicity, in the following we place our observers, Alice and Bob, at the origins of their respective spacetime diagrams with coordinates $(ct_{\rm A},x_{\rm A})$ and $(ct_{\rm B},x_{\rm B})$ where they remain at rest. In addition, we assume that Bob moves with a positive time-dependent velocity $v_{\rm B}(t_{\rm A})<c$ along a trajectory in Alice’s spacetime diagram. To answer the question, what is the quantum state $|\psi_{\rm B}\rangle$ of the EM field in the moving frame when Alice prepares the system in the state $|\psi_{\rm A}\rangle$, we notice that not only the speed of light $c$ but also its direction of propagation $s$ is invariant under reference frame transformations. This allows us to replace the spacetime coordinates $(ct_{i},x_{i})$ of the trajectories of short light pulses in each reference frame $(i={\rm A},{\rm B})$ with the natural coordinates $\chi_{i}$ where $\chi_{i}=x_{i}-sct_{i}$ [22]. Here, $s=\pm 1$ refers to light propagating in the direction of increasing and decreasing $x$ respectively. As both Alice and Bob progress along time-like trajectories, which intersect each light pulse only once, there must be a single-valued relationship between the coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ corresponding to the same light-like trajectory, but viewed from different reference frames.
@@
In general, the difference in Alice’s and Bob’s coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ is accompanied by a difference between the apparent density of the light-like trajectories in the two reference frames [22]. Even when this density differs between observers, however, the energy that each light-like worldline contributes to the EM field must be the same. To this end, Ref. [22] considers a collection of narrow light pulses seen by two inertial observers that are trapped inside a “box” co-moving with the pulses and therefore containing an invariant amount of energy. Taking into account the change in the worldline density of light and imposing energy conservation, in Ref. [22] it was possible to determine how quantum states of light transform between two inertial observers. We determined that each excitation of the EM field gained an overall factor corresponding to a Doppler factor and relating to the bunching or anti-bunching of worldlines. Nevertheless, a local excitation of the EM field remains a local excitation, and the vacuum is preserved in both reference frames. As we shall see below, the same statement holds in the case of an accelerating observer who is usually associated with the Unruh effect [1, 2, 3] (cf. Fig. 1(b)).
@@
In the remainder of this paper, we first review our local photon approach to the quantisation of light [23, 24]. In addition, we establish the relationship between Alice’s and Bob’s natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ for all possible trajectories $(ct_{\rm A},\mu_{\rm A}(t_{\rm A}))$ of the moving observer. Afterwards, we use our results to identify the transformation of the local photon annihilation and creation operators from one reference frame to another. The main purpose of this paper is to show that it is indeed possible to model the quantised EM field of a moving observer in a way that is consistent with quantum physics and with special relativity without having to replace the vacuum in one reference frame with a thermal bath in the other, as previously suggested by Unruh [3]. In addition, our theory predicts changes in the density of light-like worldlines which imply changes of the zero point energy of the quantised EM field. Similar to the moving mirrors of the Casimir effect [26, 27], an accelerating observer is therefore expected to experience forces.
@@
Local Photons. From classical electromagnetism, we know that wave packets of light can occupy definite positions in space while moving at the speed of light $c$ in a well-defined direction of propagation. Nevertheless, it has been believed for a long time that it is not possible to construct a position operator for single photons [28, 29, 30]. Moreover, in standard quantum electrodynamics, the Hamiltonian of the EM field is bounded from below, and initially localised wave packets propagate unavoidably at superluminal speeds [31, 32]. To resolve these inconsistencies and to incorporate all possible solutions of Maxwell’s equations into quantum electrodynamics, we recently derived a locally acting, unbounded field Hamiltonian to model light propagation in one dimension in position space [23, 24]. Our method requires a doubling of the standard Hilbert space of the quantised EM field [33, 34, 35, 36] and is especially useful for bringing new perceptions into the behaviour of localised photonic wave packets (cf. e.g. Refs. [22, 37, 38]).
@@
The reason for the doubling of the Hilbert space of the quantised EM field is that it allows us to introduce local carriers of light—so-called bosons localised in position (blips). Like a massive star travelling through space whose gravitational field can be felt some distance away, blips are local excitations accompanied by non-local electric and magnetic field amplitudes in their respective vicinities [39, 33, 23, 24]. Being the carriers of light, blips travel at a constant speed along straight lines in the spacetime diagram, so-called light-like trajectories with natural coordinates $\chi$. Suppose $x$ is the position of a blip with polarisation $\lambda$ and direction of propagation $s$ at time $t$ and $a^{\dagger}_{s\lambda}(x,t)$ denotes its creation operator. Then all $a^{\dagger}_{s\lambda}(x,t)$ with the same value $\chi=x-sct$ propagate along the same worldline and therefore have the same expectation value. Hence, we can replace all $a^{\dagger}_{s\lambda}(x,t)$ operators with the same natural coordinate $\chi$ by a single creation operator $a^{\dagger}_{s\lambda}(\chi)$. Linear optics experiments clearly demonstrate that photons are bosonic particles with [40]
@@
while all other commutators vanish.
Coordinate transformations. As mentioned already above, due to the relative motion of Bob with respect to Alice, their natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ are, in general, not the same, even when referring to the same light-like trajectory. However, independent of how Alice and Bob move with respect to each other, events that are localised in space and time in one reference frame are also localised in the other [22]. For example, events that occur simultaneously in Alice’s reference frame, but at different positions, occur at different times in Bob’s reference frame and vice versa. Analogously, events that occur at the same location, but at different times, in one reference frame are spatially separated in another. Moreover, we notice that the direction of propagation of light is the same in both reference frames, as illustrated in Fig. 2. In addition, we know that Alice and Bob move away from each other at the same speed, although in opposite directions. Below, we use these observations to identify the transformation between the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$.
Transformation between inertial frames. For simplicity, we first consider the scenario where Bob moves with a constant speed $v_{\rm B}>0$ away from Alice. Without restrictions, we place Alice and Bob in the following at the origin of their respective coordinate systems. Let us moreover assume that the origins of their spacetime diagrams coincides initially at $t_{\rm A}=t_{\rm B}=0$ such that Alice’s light-like trajectory with $\chi_{\rm A}=0$ corresponds to $\chi_{\rm B}=0$ in Bob’s reference frame. To establish the relationship between $\chi_{\rm A}$ and $\chi_{\rm B}$, we assume that Alice triggers an explosion at her position at a time which equals $t_{{\rm A}}^{(1)}$ and $t_{{\rm B}}^{(1)}$ in Alice’s and Bob’s reference frames respectively. The explosion sends a short light pulse to Bob, which reaches him at times $t_{{\rm A}}^{(2)}$ and $t_{{\rm B}}^{(2)}$ in Alice’s and Bob’s frames respectively. As we can see from Fig. 2, the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ that characterise the light line created by the explosion equal
@@
The above applies since the natural coordinate $\chi_{n}$ coincides with the location where its light-like worldline crosses the $x_{i}$ axis at the initial time $t_{i}=0$.
@@
with $\beta=v_{\rm B}/c$ and with $\gamma=(1-\beta^{2})^{-1/2}$. Here $s$ denotes the direction of propagation of the light travelling away from the explosion which is the same in both reference frames. See Ref. [22] for a more detailed derivation of the above variable transformation.
In special relativity, the spacetime coordinate transformations between two inertial frames is fundamental and is described by the Lorentz transformations. The above equations express the Lorentz transformations in $1+1$ dimensions in terms of the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ of a stationary and a moving observer with a constant relative velocity. It is therefore not surprising that we obtain the same result as the relativistic Doppler effect, which describes the same scenario. Suppose Alice creates little explosions that send out light from her position at regular time intervals $\Delta t_{\rm A}$. In this case, the above discussion shows that this light propagation also occurs at regular time intervals $\Delta t_{\rm B}$ in Bob’s reference frame. As we have seen above, $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ are not the same due to the relative motion between both inertial frames. The factor $\kappa$ which represents the ratio between $\Delta\chi_{\rm B}$ and $\Delta\chi_{\rm A}$ depends only on $\beta$ and on whether the moving observer travels with or against the propagation of the relevant wave packets of light.
@@
Not surprisingly, as shown in App. A, we find that the transformation between $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ for each individual line segment is nothing but the Doppler coefficient which is shown in Eq. (3). Considering the limit of small $\Delta\chi_{\rm A}$, we therefore find that Eq. (3) becomes
$\chi_{\rm B}$ $=$ $\int_{0}^{\chi_{\rm A}}{\rm d}\chi_{\rm A}^{\prime}\,\gamma(\chi_{\rm A}^{\prime})\left[1+s\beta(\chi_{\rm A}^{\prime})\right].$ (4)
in the case of a stationary and a general non-inertial observer, Alice and Bob, who are placed in the origins of their respective coordinate systems and overlap at an initial time $t_{\rm A}=t_{\rm B}=0$. In Eq. (4), the coefficients $\beta$ and $\gamma$ of non-inertial reference frames have been replaced by position-dependent functions $\beta(\chi_{\rm A}^{\prime})$ and $\gamma(\chi_{\rm A}^{\prime})$ with $\beta(\chi_{\rm A}^{\prime})=v_{\rm B}(\chi_{\rm A}^{\prime})/c$ and $\gamma(\chi_{\rm A}^{\prime})=(1-\beta(\chi_{\rm A}^{\prime})^{2})^{-1/2}$, which can be calculated analytically for any given trajectory $(ct_{\rm A}^{\prime},\mu(t_{\rm A}^{\prime}))$ of Bob with respect to Alice’s spacetime diagram, i.e. when Bob’s velocity relative to Alice is known at all of his positions $\chi_{\rm A}^{\prime}$ that he passes through before reaching his current position.
Implications for the Unruh effect. In the following, we denote the blip annihilation operators in Alice’s and in Bob’s reference frames by $a_{s\lambda}(\chi_{A})$ and $b_{s\lambda}(\chi_{B})$ respectively. As pointed out earlier, each blip operator is associated with a light-like trajectory or ray in the spacetime diagram. Since we also know that $s$ and $\lambda$ are invariant under reference frame transformations, we can fix $s$ and $\lambda$ in the following considerations without losing generality. As each light-like trajectory is common to all observers, we can now use the conservation of the blip worldlines to establish a relationship between $a_{s\lambda}(\chi_{A})$ and $b_{s\lambda}(\chi_{B})$. To do so, we consider a fixed region of space in Alice’s and Bob’s reference frames of width $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ respectively and demand that the total number, $n$, of light-like trajectories passing through is the same, even if the width of the region is not the same for both observers.
Using this invariance, we find that $n$ must equal both the expressions $\int_{\Delta\chi_{\rm B}}\text{d}\chi_{\rm B}\,b^{\dagger}_{s\lambda}(\chi_{\rm B})b_{s\lambda}(\chi_{\rm B})$ and $\int_{\Delta\chi_{\rm A}}\text{d}\chi_{\rm A}\,a^{\dagger}_{s{\sf\lambda}}(\chi_{\rm A})a_{s\lambda}(\chi_{\rm A})$. However, this equivalence is satisfied only when a local Doppler factor (cf. Eq. (4)) is added to the respective blip operator, i.e.
@@
Conclusions. In this study, we adopted a local representation of single-photon excitations [23, 24, 37] and associated each invariant light-like trajectory with blip annihilation operators $a_{s\lambda}(\chi_{\rm A})$ and $b_{s\lambda}(\chi_{\rm B})$ in the stationary and in the moving reference frames respectively. Most importantly, we established a local transformation between both sets of operators (cf. Eq. (5)) whilst avoiding the usual flat spacetime approximation. This transformation preserves the vacuum, which is, therefore, common to all observers. Our theory also predict changes in the local density of light-like worldlines in the reference frame of the moving observer which change the zero point energy of the quantised EM field, analogous to the effect of moving plates in the Casimir effect [26, 27]. Although our results are consistent with special relativity, they differ from the standard photon pair creation result, which became known as the Unruh effect [1, 2, 3]. Further theoretical investigations and additional experiments are needed to decide which theory is correct.

### v2

@@ Abstract
Whenever an experiment can be described classically, quantum physics must predict the same outcome. Intuitively, there is nothing quantum about an accelerating observer travelling through vacuum. It is, therefore, not surprising that many people are puzzled by the Unruh effect, which predicts that the observer encounters photons in a thermal state. This paper uses a recently introduced local photon approach to show that the quantised electromagnetic field in a non-inertial reference frame can be modelled without violating the principles of general relativity while both observers share a common vacuum. The only difference between a resting and an accelerating observer is that they each experience different worldline densities which implies different zero point energy densities in each reference frame.
@@
Introduction. Despite its doubtless success, quantum physics still makes some unusual predictions that continue to puzzle scientists and raise questions about the validity of some of its aspects. One example is the Unruh effect [1, 2, 3], which was proposed in 1976 and predicts that a uniformly accelerating observer will detect a thermal bath of photons where a resting observer would not see any. The creation of photon pairs in the frame of an accelerating observer seems necessary in order to reconcile relativity with quantum electrodynamics [7, 9, 4, 8, 6, 5]; however, the standard explanation of the Unruh effect does not provide much physical intuition into this process, and relies instead on mostly mathematical arguments. To confirm the Unruh effect, experiments have studied the interaction between accelerating charged particles and the surrounding free radiation field, but remained inconclusive regarding the question of photon pair creation from the vacuum [10, 11, 12, 13]. In addition, analogue gravity experiments have been performed [14, 15]. These, however, are not independent confirmations of photon pair creation in the reference frame of an accelerating observer since they are designed to exhibit a certain expected behavior.
@@
Figure 1: (a) A spacetime diagram illustrating the Doppler effect. Here, a moving observer (Bob) travels at a constant velocity $v_{\rm B}$ (green line) away from a resting observer (Alice). In addition, we assume that Alice sends light signals to Bob at intervals of $\Delta t_{\rm A}$ which he receives at constant time intervals $\Delta t_{\rm B}$. (b) A spacetime diagram illustrating the Unruh effect. Now Bob moves with a time-dependent velocity $v_{\rm B}(t_{\rm A})<c$ (green line) away from Alice. Notice that Bob’s dynamics look the same in both figures during a short time interval $\Delta t_{\rm B}$ as his trajectory can be approximated by a collection of short straight lines at velocity $v_{\rm B}^{(m)}$ with $m=1,2,3,...$
@@
An effect that is closely related to the Unruh effect, but much better understood, is the relativistic Doppler effect [16, 17, 18]. Suppose a stationary observer (let us call her Alice) sends short light pulses times $\Delta t_{\rm A}$ apart to another observer (let us call him Bob) who is moving at a constant velocity $v_{\rm B}>0$ away from Alice (cf. Fig. 1(a)). From special relativity, we know that the speed of light $c$ is the same in the reference frames of both observers, although they experience space and time differently. Since the time $\Delta t_{\rm B}$ between two subsequent light pulses arriving at Bob’s position is larger than $\Delta t_{\rm A}$, time passes more slowly in the moving frame. Another way of describing the Doppler effect is to notice that the frequency of monochromatic waves of light is lower in Bob’s reference frame than in Alice’s. Most importantly, if the electromagnetic (EM) field is in its vacuum state in one frame, the same is true in the other. Taking this into account, the particle generation predicted by the Unruh effect seems even more mysterious since the trajectory of an accelerating observer can be approximated by a collection of short time intervals during which the observer moves at a constant speed $v_{\rm B}^{(m)}$ with $m=1,2,3,...$ (cf. Fig. 1(b)).
@@
In this paper, we therefore revisit the validity of the standard conclusions of light propagation in non-inertial reference frames by making use of a recent physically motivated quantisation of the EM field in position space [19, 20]. The main advantage of considering local instead of monochromatic photons as the basic building blocks of the quantised EM field is that this allows us to make use of the principle of relativity and of spatial and time translational symmetries in a more direct way. Moreover, we can avoid the approximation of standing modes in a flat spacetime, which usually accompanies the derivation of the Unruh effect [21]. As we shall see below, our alternative description of the EM field suggests that the vacuum state is the same in all non-inertial reference frames —only zero point energies change. Consequently, the situation in Fig. 1(b) can be modelled consistently without the need for thermal photons.
@@
For simplicity, in the following we place our observers, Alice and Bob, at the origins of their respective spacetime diagrams, with coordinates $(ct_{\rm A},x_{\rm A})$ and $(ct_{\rm B},x_{\rm B})$, where they remain at rest. In addition, we assume that Bob moves with a positive time-dependent velocity $v_{\rm B}(t_{\rm A})<c$ along a trajectory in Alice’s spacetime diagram. To answer the question, what is the quantum state $|\psi_{\rm B}\rangle$ of the EM field in the moving frame when Alice prepares the system in the state $|\psi_{\rm A}\rangle$, we notice that not only the speed of light $c$, but also its direction of propagation $s$ is invariant under reference frame transformations. This allows us to replace the spacetime coordinates $(ct_{i},x_{i})$ of the trajectories of short light pulses in each reference frame $(i={\rm A},{\rm B})$ with the natural coordinates $\chi_{i}$ where $\chi_{i}=x_{i}-sct_{i}$ [18], which are unique to each light pulse. Here, $s=\pm 1$ refers to light propagating in the direction of increasing and decreasing $x$ respectively. As both Alice and Bob progress along time-like trajectories, which intersect each light pulse only once, there must be a single-valued relationship between the coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ corresponding to the same light-like trajectory, but viewed from different reference frames.
@@
In general, the difference in Alice’s and Bob’s coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ is accompanied by a difference between the apparent density of the light-like trajectories in the two reference frames [18]. Even when this density differs between observers, however, the energy that each light-like worldline contributes to the EM field must be the same. For this reason, in Ref. [18] we considered a collection of narrow light pulses seen by two inertial observers trapped inside a “box” co-moving with the pulses, which therefore contained an invariant amount of energy. Taking into account the change in the worldline density of light and imposing energy conservation, in Ref. [18] it was then possible to determine how quantum states of light transform between two inertial observers. We determined that each excitation of the EM field gained an overall factor corresponding to a Doppler factor and relating to the bunching or anti-bunching of worldlines. Nevertheless, a local excitation of the EM field remains a local excitation, and the vacuum is preserved in both reference frames. As we shall see below, the same statement holds in the case of an accelerating observer who is usually associated with the Unruh effect [1, 2, 3] (cf. Fig. 1(b)).
@@
In the remainder of this paper, we first review our local photon approach to the quantisation of light [19, 20]. In addition, we establish the relationship between Alice’s and Bob’s natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ for all possible trajectories $(ct_{\rm A},\mu_{\rm A}(t_{\rm A}))$ of the moving observer. Afterwards, we use our results to identify the transformation of the local photon annihilation and creation operators from one reference frame to the other. The main purpose of this paper is to show that it is indeed possible to model the quantised EM field of a moving observer in a way that is consistent with quantum physics and with special relativity without having to replace the vacuum in one reference frame with a thermal bath in the other, as previously suggested by Unruh [3]. In addition, our theory predicts changes in the density of light-like worldlines which imply changes of the zero point energy of the quantised EM field. Similar to the moving mirrors of the Casimir effect [22, 23], an accelerating observer is therefore expected to experience forces and energy shifts.
@@
Local Photons. From classical electromagnetism, we know that wave packets of light can occupy definite positions in space while moving at the speed of light $c$ in a well-defined direction of propagation. To incorporate all possible solutions of Maxwell’s equations into quantum electrodynamics, we recently introduced a local photon model to describe light propagation along the $x$ axis [19, 20]. Our approach distinguishes between monochromatic photons with positive and negative frequencies [24, 25, 26, 27] and is especially useful for bringing new perceptions into the behaviour of localised photonic wave packets (cf. e.g. Refs. [18, 28, 29, 23]). After doubling the standard Hilbert space of the quantised EM field, we could overcome previous objections regarding the existence of a position operator for single photons [30, 31, 32]. In addition, we obtained a locally acting, unbounded field Hamiltonian such that initially localised wave packets no longer travel at superluminal speeds, as it is the case for bounded Hamiltonians [33, 34].
@@
In our local photon approach, we introduce local carriers of light—so-called bosons localised in position (blips)—which offer an alternative basis for the construction of single-photon wave packets. In agreement with general relativity, blips travel at a constant speed along straight lines in the spacetime diagram, so-called light-like trajectories with natural coordinates $\chi$. Suppose $x$ is the position of a blip with polarisation $\lambda$ and direction of propagation $s$ at time $t$ and $a^{\dagger}_{s\lambda}(x,t)$ denotes its creation operator. Then all $a^{\dagger}_{s\lambda}(x,t)$ with the same value $\chi=x-sct$ propagate along the same worldline and therefore have the same expectation value. Hence, we can replace all $a^{\dagger}_{s\lambda}(x,t)$ operators with the same natural coordinate $\chi$ by a single creation operator $a^{\dagger}_{s\lambda}(\chi)$. Linear optics experiments clearly demonstrate that photons are bosonic particles with [35]
@@
while all other commutators vanish. Nevertheless, similar to how a massive star carries a gravitational field that can be felt some distance away, blips are local carriers of non-local electric and magnetic field amplitudes [36, 24, 19, 20]. For example, the electromagnetic field associated with the coherent state of a blip annihilation operator $a_{s\lambda}(\chi)$ can in principle be felt anywhere along the $x$ axis, but is strongest nearest the blip itself.
Coordinate transformations. As mentioned already above, due to the relative motion of Alice and Bob, their natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ are, in general, not the same, even when referring to the same light-like trajectory. However, independent of how Alice and Bob move with respect to each other, events that are localised in space and time in one reference frame are also localised in the other [18]. For example, events that occur simultaneously in Alice’s reference frame, but at different positions, occur at different times in Bob’s reference frame and vice versa. Analogously, events that occur at the same location, but at different times, in one reference frame are spatially separated in another. Moreover, we notice that the direction of propagation of light is the same in both reference frames, as illustrated in Fig. 2. In addition, we know that Alice and Bob move away from each other at the same speed, although in opposite directions. Below, we use these observations to identify the transformation between the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$.
Transformation between inertial frames. For simplicity, we first consider the scenario where Bob moves with a constant speed $v_{\rm B}>0$ away from Alice. Without restrictions, we place Alice and Bob in the following at the origins of their respective coordinate systems. Let us moreover assume that the origins of their spacetime diagrams coincide initially at $t_{\rm A}=t_{\rm B}=0$ such that Alice’s light-like trajectory with $\chi_{\rm A}=0$ corresponds to $\chi_{\rm B}=0$ in Bob’s reference frame. To establish the relationship between $\chi_{\rm A}$ and $\chi_{\rm B}$, we assume that Alice triggers an explosion at her position at a time which equals $t_{{\rm A}}^{(1)}$ and $t_{{\rm B}}^{(1)}$ in Alice’s and Bob’s reference frames respectively. The explosion sends a short light pulse to Bob, which reaches him at a time that equals $t_{{\rm A}}^{(2)}$ and $t_{{\rm B}}^{(2)}$ in Alice’s and Bob’s frames respectively. As we can see from Fig. 2, the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ that characterise the light line created by the explosion equal
@@
The above applies since the natural coordinate $\chi_{n}$ coincides with the location where the light-like worldline crosses the $x_{i}$ axis at the initial time $t_{i}=0$.
@@
with $\beta=v_{\rm B}/c$ and with $\gamma=(1-\beta^{2})^{-1/2}$. Here $s$ denotes the direction of propagation of the light travelling away from the explosion, which is the same in both reference frames. See Ref. [18] for a more detailed derivation of the above variable transformation.
In special relativity, the spacetime coordinate transformations between two inertial frames is fundamental and is described by the Lorentz transformations. The above equations express the Lorentz transformations in $1+1$ dimensions in terms of the natural coordinates $\chi_{\rm A}$ and $\chi_{\rm B}$ of a stationary and a moving observer with a constant relative velocity. It is therefore not surprising that we obtain the same result as the relativistic Doppler effect, which describes the same scenario. Suppose Alice creates little explosions that send out light from her position at regular time intervals $\Delta t_{\rm A}$. In this case, the above discussion shows that the light pulses are received at regular time intervals $\Delta t_{\rm B}$ in Bob’s reference frame. As we have seen above, $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ are not the same due to the relative motion between both inertial frames. The factor $\kappa$ which represents the ratio between $\Delta\chi_{\rm B}$ and $\Delta\chi_{\rm A}$ depends only on $\beta$ and on whether the moving observer travels with or against the propagation of the relevant wave packets of light.
@@
Not surprisingly, as shown in App. A, we find that the transformation between $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ for each individual line segment is nothing but the Doppler coefficient which is given in Eq. (3). Considering the limit of small $\Delta\chi_{\rm A}$, we therefore find that Eq. (3) becomes
$\chi_{\rm B}$ $=$ $\int_{0}^{\chi_{\rm A}}{\rm d}\chi_{\rm A}^{\prime}\,\gamma(\chi_{\rm A}^{\prime})\left[1+s\beta(\chi_{\rm A}^{\prime})\right]$ (4)
in the case of a stationary and a general non-inertial observer, Alice and Bob, who are placed at the origins of their respective coordinate systems and overlap at an initial time $t_{\rm A}=t_{\rm B}=0$. In Eq. (4), the coefficients $\beta$ and $\gamma$ of non-inertial reference frames have been replaced by position-dependent functions $\beta(\chi_{\rm A}^{\prime})$ and $\gamma(\chi_{\rm A}^{\prime})$ with $\beta(\chi_{\rm A}^{\prime})=v_{\rm B}(\chi_{\rm A}^{\prime})/c$ and $\gamma(\chi_{\rm A}^{\prime})=(1-\beta(\chi_{\rm A}^{\prime})^{2})^{-1/2}$, which can be calculated analytically for any given trajectory $(ct_{\rm A}^{\prime},\mu(t_{\rm A}^{\prime}))$ of Bob with respect to Alice’s spacetime diagram, i.e. when Bob’s velocity relative to Alice is known at all positions $\chi_{\rm A}^{\prime}$ that he passes through before reaching his current position.
Implications for the Unruh effect. In the following, we denote the blip annihilation operators in Alice’s and Bob’s reference frames by $a_{s\lambda}(\chi_{A})$ and $b_{s\lambda}(\chi_{B})$ respectively. As pointed out earlier, each blip operator is associated with a light-like trajectory or ray in the spacetime diagram. Since we also know that $s$ and $\lambda$ are invariant under reference frame transformations, we can fix $s$ and $\lambda$ in the following considerations without losing generality. As each light-like trajectory is common to all observers, we can now use the conservation of the blip worldlines to establish a relationship between $a_{s\lambda}(\chi_{A})$ and $b_{s\lambda}(\chi_{B})$. To do so, we consider a fixed region of space in Alice’s and Bob’s reference frames of width $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}$ respectively and demand that the total number, $n$, of light-like trajectories passing through this region is the same, even if the width of the region is not the same for both observers.
Using this invariance, we find that $n$ must equal both $\int_{\Delta\chi_{\rm B}}\text{d}\chi_{\rm B}\,b^{\dagger}_{s\lambda}(\chi_{\rm B})b_{s\lambda}(\chi_{\rm B})$ and $\int_{\Delta\chi_{\rm A}}\text{d}\chi_{\rm A}\,a^{\dagger}_{s{\sf\lambda}}(\chi_{\rm A})a_{s\lambda}(\chi_{\rm A})$. However, this equivalence is satisfied only when a local Doppler factor (cf. Eq. (4)) is added to the respective blip operator, i.e.
@@
Conclusions. In this study, we adopted a local representation of single-photon excitations [19, 20, 29] and associated each invariant light-like trajectory with blip annihilation operators $a_{s\lambda}(\chi_{\rm A})$ and $b_{s\lambda}(\chi_{\rm B})$ in the stationary and in the moving reference frames respectively. The local transformation between both sets of operators (cf. Eq. (5)) can be established whilst avoiding the flat spacetime approximation which is usually made in derivations of the Unruh effect [21]. Our theory predicts changes in the local density of light-like worldlines in the reference frame of the moving observer which alters the zero point energy of the quantised EM field. Nevertheless, the vacuum remains common to all observers. In this way, our observations differ from the standard photon pair creation result which became known as the Unruh effect [1, 2, 3]. Our approach is therefore likely to stimulate further research into quantum electrodynamics in non-inertial reference frames.